# Full-scale Representation Guided Network for Retinal Vessel Segmentation

### v1

@@ Abstract
The U-Net architecture and its variants have remained state-of-the-art (SOTA) for retinal vessel segmentation over the past decade. In this study, we introduce a Full Scale Guided Network (FSG-Net), where the feature representation network with modernized convolution blocks extracts full-scale information and the guided convolution block refines that information. Attention-guided filter is introduced to the guided convolution block under the interpretation that the filter behaves like the unsharp mask filter. Passing full-scale information to the attention block allows for the generation of improved attention maps, which are then passed to the attention-guided filter, resulting in performance enhancement of the segmentation network. The structure preceding the guided convolution block can be replaced by any U-Net variant, which enhances the scalability of the proposed approach. For a fair comparison, we re-implemented recent studies available in public repositories to evaluate their scalability and reproducibility. Our experiments also show that the proposed network demonstrates competitive results compared to current SOTA models on various public datasets. Ablation studies demonstrate that the proposed model is competitive with much smaller parameter sizes. Lastly, by applying the proposed model to facial wrinkle segmentation, we confirmed the potential for scalability to similar tasks in other domains. Our code is available on https://github.com/ZombaSY/FSG-Net-pytorch.
1 Introduction
@@
Convolutional neural networks (CNNs) have seen significant improvements in performance and optimization since the 2010s. The introduction of hardware acceleration using GPUs, ReLU activation function, and residual block [11, 1] has enabled smooth back-propagation in deep neural network architectures. Research in this field has focused on finding a balance between computational efficiency, parameter size, code scalability, and fidelity. Depthwise separable convolution [5] and squeeze-and-excitation [12] have been particularly influential in this regard. The inverted residual block [19] achieved higher fidelity with optimized computational efficiency and a smaller parameter size compared to ResNet.
@@
Figure 1: F1 scores of compared networks on the DRIVE dataset, measured against the validation dataset comprising zero-padded images of resolution 608$\times$608. Among the considered architectures, FSG-Net-T achieved a superior F1 score compared to competitive models while maintaining a reduced parameter size relative to its counterparts. Additionally, the FSG-Net-L achieved the highest F1 scores while possessing a median parameter size.
@@
In the evolutionary history of CNNs, a noteworthy highlight is the dominance of U-Net [18] and its variants [14, 21, 20, 28, 13, 8, 22] as SOTA models in the field of medical image segmentation, especially for retinal vessel segmentation tasks on well-known datasets such as DRIVE, STARE, CHASE_DB1, and HRF. On the contrary, for clinical segmentation tasks on the CVC-clinic and Kvasir-SEG datasets, vision transformer models are employed [23, 2, 3, 9]. The primary distinction between retinal vessels and other clinical datasets lies in the level of feature intricacy required. Vision transform models exhibit some limitations in overcoming their constrained inductive bias. Meanwhile, our experiments underscore the ongoing relevance and effectiveness of attention mechanisms in addressing such challenges.
@@
In this study, we aimed to propose a U-Net-based segmentation network that reflects the thin and elongated structural characteristics of retinal vessels, starting with the lower layers and gradually working our way up to the upper layers. Furthermore, based on the understanding that the guided filter [10] can function as a type of edge sharpening filter, the guided filter is located at the decoder after features from the encoder are merged so that full scale features can be used as input to the guided filters at each stage. Our network architecture improves both performance and computational efficiency simultaneously as shown in Figure 1. To ensure a fair comparison of the robustness of competing models, the training environment was maintained consistently.
The main contributions of this paper are summarized as follows: Firstly, enhancement of feature extraction through the proposal of a novel and efficient convolutional block. Secondly, proposing guided convolution blocks in the decoder, demonstrating performance enhancement as a result and enabling integration with various U-Net segmentation models. Thirdly, a comparison of various SOTA algorithms under a fixed experimental environment.
2 Method
2.1 Motivation
@@
$\hat{I_{i}}=a_{k}I_{i}+b_{k},\quad\forall i\in w_{k},$ (1)
where $a$, $b$ are linear coefficients, and $w$ is a local window. Here, we show why the guided filter can improve edge-like blood vessel segmentation performance. If we consider overlapping windows, the $i$-th pixel has several $a_{k}$ and $b_{k}$ values depending on the window size. Thus, the guided filter output of $I_{i}$ can be averaged as follows:
$\hat{I_{i}}=\frac{1}{|w|}\sum_{k\in{w_{i}}}\left(a_{k}I_{i}+b_{k}\right),\quad\forall i\in w_{k}.$ (2)
@@
where ()¯$\bar{\left(\cdot\right)}_{k}$ means average in $w_{k}$. Then, by putting Eq. (3) into Eq. (2), we have the following formulation:
$\hat{I}_{i}=\bar{a}_{i}(I_{i}-\bar{I}_{i})+\tilde{p}_{i},$ (4)
@@
where ()~$\tilde{\left(\cdot\right)}_{k}$ denotes the average of the average, meaning $\tilde{p}_{i}=\frac{1}{|w|}\sum_{k\in w_{i}}{\bar{p}}_{k}$. The unsharp mask filtering is defined as $\hat{I}=\alpha\left(I-L(I)\right)+I$, where $L(\cdot)$ is a low-pass filter such as the Gaussian filter. Eq. (4) looks like unsharping mask filter, where sharpening mask $\left(I_{i}-\bar{I}_{i}\right)$ from the guidance image is added to the averaged target image. The filter strength is controlled by $\alpha$.
@@
In their pioneering work [26], Zhang et al. adopted the guided filter in the segmentation network and suggested incorporating the attention map $M$ into the minimization problem to estimate $a$ and $b$:
$E(a_{k},b_{k})=\sum_{i\in{w_{k}}}\left(M_{i}^{2}\left(a_{k}{I_{d}}_{i}+b_{k}-g_{i}\right)^{2}+\mu a_{k}^{2}\right).$ (5)
@@
In Eq. (5), the ${I_{d}}_{i}$ is the downsampled input feature from encoding parts, and the $g_{i}$ is the gating signal from upward path at the decoding parts. Eq. (5) implies that an improved attention map can lead to better solutions. In the following, we are going to propose a method that can generate an improved attention map with full-scale information.
2.2 Network architecture
Figure 2: Network architecture of the proposed FSG-Net.
@@ 2.2 Network architecture
As can be seen in Figure 2, the proposed network basically follows the U-Net architecture but has an additional guided convolution block after the feature representation network, unlike usual U-Net based structures. Contrary to the standard U-Net consisting of five stages, we have opted for four stages in FSG-Net based on the assumption that a wider receptive field is less critical for retinal vessel segmentation.
@@
As described in the left part of FSG-Net, which is called the feature representation network, the input layer of the down-convolution is concatenated with a separate convolution from the other layers. These preserved features are then connected to the up-convolution layer. The deep bottleneck structure of the convolution block, indicated by the thick arrows, delivers a more enhanced feature representation. In the feature representation network, the key area to note is the feature merging section, indicated by the red dashed box. In this feature merging process, features from three paths—the current, upper, and lower stages—are concatenated and passed through a newly designed convolution block before being forwarded to the subsequent stages. This approach allows the integration and transmission of information across all scales.
@@
In the right part of the model, namely the guided convolution block, compressed features reflecting full scale information are given as the current stage input to the guided residual module (GRM). Up-stage input from the higher stage is given as another input for guided filtering. After GRM, predictions are derived through convolution and activation. For training, a combination of the dice loss and BCE loss was employed to improve segmentation performance of the edge-like structures [7]:
$\begin{split}L_{\text{DS}}=\sum_{d=1}^{D}\sum_{i=0}^{S^{d}}\left(y_{i}\log\hat{y}_{i}^{d}+(1-y_{i})\log(1-\hat{y}_{i}^{d})\right),\end{split}$ (6)
$\begin{split}L_{Dice}=1-\sum_{d=1}^{D}\sum_{i=1}^{S^{d}}\frac{2\left(y_{i}\cdot\hat{y}_{i}^{(d)}\right)}{y_{i}+\hat{y}_{i}^{(d)}+\epsilon},\end{split}$ (7)
$\begin{split}L_{total}=L_{DS}+L_{Dice},\end{split}$ (8)
where $S$ denotes the indices and $L_{Dice}$ for Dice loss over pixels. For deep-supervision, $D$ represents the number of layers, the prediction at layer $d$, and weights for each layer, respectively.
@@
2.3 Modernized convolution block
The standard U-Net has a fundamental structure comprising skip connections and double convolution blocks. Recent studies have shown that incorporating different CNN structures can significantly improve the performance of the original model [16, 9]. Inspired by the ConvNeXt [16], we have designed a convolutional block suitable for retinal vessel segmentation.
@@ 2.3 Modernized convolution block
Figure 3(a) and (b) represent the depthwise residual block and inverted residual block, respectively. Figure 3(c) shows the proposed convolution block. The structure of the block is an extension of the latest advancements in convolutional block development, characterized by features such as 1$\times$1 convolution, inverted bottleneck, and depth-wise convolution. Like other modernized convolution blocks, the proposed block employs a convolution with kernel size 2 and stride 2 in the first stage. By incorporating spatial and dimensional changes within the first block, we enable the design of a deep bottleneck stage in the down-convolution structure. This approach not only increases the information on the feature but also separates it from the bottleneck, allowing for more detailed feature representation. Furthermore, to maintain linearity, we utilize unique ReLU activation between each inverted residual block. For regularization purposes, we employ a learnable gamma parameter and apply a drop block before joining the identity block.
2.4 Guided convolution block
As shown in Figure 2, the guided convolution block consists of GRM to refine the input from the feature representation network and a convolution to output the prediction map. Figure 4 shows the overall process of GRM. The purpose of GRM is to refine the feature from the feature representation network at the same stage. In GRM, both the features at the current stage and at the higher stage are used to generate an attention map, and then the attention map is multiplied with the current feature to generate the output, which is described in the right part of Figure 4. Here we can expect an improved attention map $M$ can be estimated to solve Eq. (5) because the input feature from the feature representation network has full scale information. To further enhance the feature refinement, a residual block is introduced after the attention-guided filtering, as shown in the left part of Figure 4. By passing the features through the residual block, a stable map is generated for both the deep supervision and subsequent layers. Moreover, we incorporate a $1\times 1$ convolution to preserve the semantic information present in the feature maps.
3 Experiments
3.1 Training Techniques
In retinal vessel segmentation, the performance of studies is often determined by subtle gaps. We believe that these subtle gaps are highly influenced by the choice of hyperparameters and training/inference environment. To address this imbalance, we fixed all hyperparameters for training and inference. We empirically found that RandAugment [6] with a specific scale did not work well on medical datasets; therefore, we customized it to better suit our datasets. Training techniques include blur, color jitter, horizontal flip, perspective transformation, resize, crop, and CutMix [25].
@@ 3.1 Training Techniques
lr scheduler Cosine annealing
@@
random blur prob=0.8
random jitter prob=0.8
@@
random perspective prob=0.3
random random resize prob=0.8
@@
3.2 Implementation Details
Our experimental environment comprises an Intel Xeon Gold 5220 processor, a Tesla V100-SXM2-32GB GPU, Pytorch 1.13.1, and CUDA version 11.7. The inference time for FSG-Net-L was approximately $600ms$ for an input size of (608, 608) using a GPU-synchronized flow. To address as much variability as possible, we re-implemented comparison studies and integrated them into a single environment. To ensure experimental fairness, certain hyperparameters, including the framework, loss function, metric, data augmentation, and random seed, were fixed to measure the robustness of the model. Our training settings followed the hyperparameters in Table  1, used for segmentation tasks in ADE20K multiscale learning in ConvNeXt [16].
@@ 3.2 Implementation Details
3.3 Datasets
The DRIVE dataset comprised 40 retinal images with a resolution of 565$\times$584 pixels, captured as part of a retinopathy screening study in the Netherlands. The STARE dataset comprises 20 retinal fundus images with a resolution of 700 × 605 pixels, and the CHASE_DB1 dataset includes 28 retinal images from schoolchildren with a resolution of 999$\times$960 pixels. Both the STARE and CHASE_DB1 datasets were manually annotated by two independent experts. We used the annotation of the first expert, named "Hoover A." in STARE and "1stHO" in CHASE_DB1, for our analysis. The HRF dataset comprises 45 images, equally divided into a 1:1:1 ratio of healthy patients, diabetic retinopaths, and glaucomatous patients, with a high resolution of 3504$\times$2336 pixels. To measure the performance of the models, it is necessary to divide the data into training and validation sets. As the retinal vessel segmentation dataset was relatively limited, we split the data into a 1:1 ratio of the training and validation sets. The DRIVE dataset was officially divided into training and validation sets, each containing 20 images. For the STARE, CHASE_DB1, and HRF datasets, we used the first half as training and the remaining half as valiation.
Table 2: Metric comparison of results on datasets.
  Architecture mIoU F1 score Acc AUC Sen MCC Rank Avg
@@ 3.3 Datasets
U-Net 83.857 82.956 97.013 97.853 83.449 81.456 5
U-Net++ 81.228 79.564 96.524 96.271 77.802 77.830 13
U-Net3+ Deep 83.909 83.030 97.017 98.082 83.721 81.520 3.1
ResU-Net 83.862 82.953 97.021 97.766 83.226 81.453 5.8
ResU-Net++ 83.729 82.783 97.001 97.708 82.791 81.263 9.1
SAU-Net 83.368 82.334 96.925 97.616 82.311 80.782 11.2
DCASU-Net 83.743 82.808 96.996 97.838 83.080 81.290 7.8
AG-Net 83.176 82.111 96.882 97.628 82.155 80.540 11.8
AttU-Net 83.958 83.080 97.039 97.844 83.422 81.584 3.2
R2U-Net 83.555 82.580 96.952 97.879 82.961 81.038 8.8
ConvU-NeXt 83.800 82.882 97.012 97.835 83.019 81.367 7.2
FR-UNet 83.884 82.995 97.007 98.158 83.869 81.485 3.8
FSG-Net-L (ours) 84.068 83.229 97.042 98.235 84.207 81.731 1
@@
U-Net 85.924 84.873 97.754 98.341 84.361 83.713 3.3
U-Net++ 81.514 79.061 97.022 94.479 75.539 77.764 12.8
U-Net3+ Deep 85.824 84.829 97.707 99.146 85.522 83.626 3.5
ResU-Net 85.964 84.872 97.767 98.050 83.997 83.726 3.5
ResU-Net++ 83.185 81.358 97.319 95.221 78.456 80.196 11.7
SAU-Net 85.158 84.061 97.604 97.784 84.015 82.885 7
DCASU-Net 84.423 83.064 97.454 98.362 83.583 81.771 8.8
AG-Net 84.811 83.766 97.565 98.403 83.347 82.516 7.7
AttU-Net 85.848 84.772 97.694 99.050 86.226 83.588 3.8
R2U-Net 83.727 81.786 97.468 93.457 77.048 80.810 11.3
ConvU-NeXt 85.339 84.186 97.658 97.866 83.401 82.998 6.7
FR-UNet 84.815 83.496 97.577 96.637 81.872 82.327 9.2
FSG-Net-L (ours) 86.118 85.100 97.746 98.967 86.608 83.958 1.7
@@
U-Net 82.065 80.159 97.404 99.368 85.370 79.002 6
U-Net++ 81.512 79.415 97.321 99.362 84.143 78.201 8.2
U-Net3+ Deep 82.489 80.697 97.483 99.506 85.740 79.558 2.7
ResU-Net 81.104 78.810 97.320 99.278 81.311 77.479 11.5
ResU-Net++ 73.276 67.411 95.966 96.223 67.782 65.593 13
SAU-Net 81.335 79.100 97.338 99.407 82.265 77.802 8.8
DCASU-Net 82.254 80.368 97.483 99.329 83.916 79.161 5.3
AG-Net 82.158 80.272 97.440 99.544 84.817 79.070 4.2
AttU-Net 82.562 80.742 97.546 99.430 83.907 79.537 3.3
R2U-Net 81.250 78.944 97.384 98.792 80.054 77.625 10.8
ConvU-NeXt 81.752 79.704 97.405 99.402 83.009 78.439 7
FR-UNet 81.330 79.170 97.269 99.544 84.744 77.944 7.7
FSG-Net-L (ours) 82.680 81.019 97.515 99.378 85.995 79.899 2.2

### v2

@@ Abstract
The U-Net architecture and its variants have remained state-of-the-art (SOTA) for retinal vessel segmentation over the past decade. In this study, we introduce a Full-Scale Guided Network (FSG-Net), where a novel feature representation module using modernized convolution blocks effectively captures full-scale structural information, while a guided convolution block subsequently refines this information. Specifically, we introduce an attention-guided filter within the guided convolution block, leveraging its similarity to unsharp masking to enhance fine vascular structures. Passing full-scale information to the attention block facilitates the generation of more contextually relevant attention maps, which are then passed to the attention-guided filter, providing further refinement to the segmentation performance. The structure preceding the guided convolution block can be replaced by any U-Net variant, ensuring flexibility and scalability across various segmentation tasks. For a fair comparison, we re-implemented recent studies available in public repositories to evaluate their scalability and reproducibility. Our experiments demonstrate that, despite its compact architecture, FSG-Net delivers performance competitive with SOTA methods across multiple public datasets. Ablation studies further demonstrate that each proposed component meaningfully contributes to this competitive performance. Our code is available on https://github.com/ZombaSY/FSG-Net-pytorch.
Introduction
@@
Convolutional neural networks (CNNs) have seen significant improvements in performance and optimization since the 2010s. The introduction of hardware acceleration using GPUs, ReLU activation function, and residual block [He_2016_CVPR, https://doi.org/10.48550/arxiv.1802.06955] has enabled smooth back-propagation in deep neural network architectures. Research in this field has focused on finding a balance between computational efficiency, parameter size, code scalability, and fidelity. Depthwise separable convolution [Chollet_2017_CVPR] and squeeze-and-excitation [Hu_2018_CVPR] have been particularly influential in this regard. The inverted residual block [Sandler_2018_CVPR] achieved higher fidelity with optimized computational efficiency and a smaller parameter size compared to ResNet.
@@
Figure 1: F1 scores of compared networks on the DRIVE dataset, measured against the validation dataset comprising zero-padded images of resolution $\times608$. Among the considered architectures, FSG-Net-T achieved a superior F1 score compared to competitive models while maintaining a reduced parameter size relative to its counterparts. Additionally, the FSG-Net achieved the highest F1 scores while possessing a median parameter size.
In the evolutionary history of CNNs, a noteworthy highlight is the dominance of U-Net [10.1007/978-3-319-24574-4_28] and its variants as SOTA models in the field of medical image segmentation across a wide spectrum of segmentation tasks [8959021, Tang_2022_CVPR, 9446143, 10.1007/978-3-030-00889-5_1, 9053405, 9413346, TIAN2020117, jin2022semi, jin2024inter, jin2021cascade, jin2021domain, tan2024deep, ding2024rcar]. In particular, retinal-vessel segmentation benefits from well-curated public datasets such as DRIVE, STARE, CHASE_DB1, and HRF, on which the latest CNN-based methods are already systematically benchmarked and firmly established as strong baselines. In [qin2024review], traditional and the recent deep learning based algorithms for retinal vessel segmentation are comprehensively reviewed.
@@
On the contrary, for clinical segmentation tasks on the CVC-clinic and Kvasir-SEG datasets, vision transformer (ViT) models are employed [xie2021segformer, 10.1007/978-3-031-25066-8_9, chen2021transunet, HAN2022109512]. The primary distinction between retinal vessels and other clinical datasets lies in the level of feature intricacy required. ViT models exhibit some limitations in overcoming their constrained inductive bias. Meanwhile, our experiments underscore the ongoing relevance and effectiveness of attention mechanisms in addressing such challenges.
@@
In this study, we aimed to propose a U-Net-based segmentation network that explicitly captures and preserves the thin and elongated structural characteristics of retinal vessels by leveraging a novel, full-scale feature representation strategy from the early layers upward. Furthermore, we introduced an attention-guided filtering approach inspired by guided filters [6319316], strategically embedded in the decoder following encoder-decoder feature merging. This design allows the guided filter to perform adaptive edge enhancement, specifically utilizing comprehensive, full-scale structural information at each decoding stage. Our network architecture improves both performance and computational efficiency simultaneously, as shown in Fig. 1. To ensure a fair comparison of the robustness of competing models, the training environment was maintained consistently.
The main contributions of this paper are summarized as follows: Firstly, a novel convolutional block specifically designed for retinal vessel segmentation, capable of extracting comprehensive, full-scale structural features with computational efficiency. Secondly, an attention-guided convolutional block integrated into the decoder, enhancing segmentation accuracy through adaptive refinement, and offering flexibility to be combined seamlessly with various U-Net architectures. Thirdly, a thorough comparison and reproducibility study of multiple SOTA algorithms, conducted under strictly controlled experimental conditions.
Related work
Guided filter
@@
$\hat{I_{i}}=a_{k}I_{i}+b_{k},\quad\forall i\in\mathrm{w}_{k},$ (1)
where $a_{k}$, $b_{k}$ are linear coefficients and assumed to be constants in the local window $\mathrm{w}_{k}$. $\hat{I}_{i}$ is the output of the guided filter with input $I_{i}$. Here, we show why the guided filter can improve edge-like blood vessel segmentation performance. If we consider overlapping windows, the $i$-th pixel has several $a_{k}$ and $b_{k}$ values depending on the window size. Thus, the guided filter output of $I_{i}$ can be averaged as follows:
$\hat{I_{i}}=\frac{1}{|w|}\sum_{k\in{w_{i}}}\left(a_{k}I_{i}+b_{k}\right),\quad\forall i\in\mathrm{w}_{k}.$ (2)
The guided filter smooths images while preserving edge information, making it widely used in various image processing tasks such as image detail enhancement, image denoising, image preprocessing, and image dehazing.
Multiscale architectures preserving spatial detail
In this study, we focus on two representative types of multiscale architectures for image segmentation: full-resolution preserving and full-scale aggregation approaches. In the full-resolution preserving category, HRNet maintains high-resolution representations by processing multi-resolution branches in parallel throughout the network [Sun_2019_CVPR]. Similarly, FR-UNet adopts a parallel full-resolution pathway while preserving the U-Net structure [9815506]. In the full-scale aggregation category, UNet3+ extends the skip-connection scheme of the original U-Net by enabling each decoder stage to simultaneously receive and fuse feature maps from all encoder and decoder stages [9053405].
Deep supervision
Deep supervision has recently become a prevalent strategy in various segmentation tasks, particularly within medical image analysis, to enhance model performance [10.1007/978-3-030-00889-5_1, wang2019deeply, reiss2021every]. In general, deep supervision refers to computing multiple loss functions at different intermediate layers of the neural network, enabling more stable gradient propagation throughout the training process. This approach has proven beneficial in reducing issues such as vanishing gradients, thus improving segmentation accuracy.
Several well-known variations of U-Net architectures, including M-Net [fu2018joint], AG-Net [10.1007/978-3-030-32239-7_88], and U-Net3+ [9053405], have successfully applied deep supervision and demonstrated better segmentation results compared to architectures utilizing a single loss function. Specifically, for retinal vessel segmentation, M-Net and AG-Net calculated multiple losses from each of the decoders, except the deepest one. On the other hand, U-Net3+ utilized losses from all decoder stages and employed them to train a segmentation model targeting liver and spleen structures in CT images.
Method
Motivation
@@
where $\bar{\left(\cdot\right)}_{k}$ means average in a window $\mathrm{w}_{k}$ and $\mu_{k}$ is the mean of $I_{k}$. Then, by putting Eq. (3) into Eq. (2), we have the following formulation:
$\hat{I}_{i}=\bar{a}_{i}(I_{i}-\mu_{i})+\tilde{p}_{i},$ (4)
where $\tilde{\left(\cdot\right)}_{k}$ denotes the average of the average, meaning $\tilde{p}_{i}=\frac{1}{|w|}\sum_{k\in w_{i}}{\bar{p}}_{k}$. The unsharp masking is defined as follows:
$\hat{I}_{i}=\alpha\left(I_{i}-\sum_{k\in\mathrm{w}_{i}}w_{ik}I_{k}\right)+I_{i},$ (5)
@@
where $w_{ik}$ denotes the normalized Gaussian weight centered at pixel $i$, evaluated at neighboring pixel $k$. Eq. (4) looks like Eq. (5), where sharpening mask $\left(I_{i}-\bar{I}_{i}\right)$ from the guidance image is added to the averaged target image. The filter strength is controlled by $\alpha$.
@@
In their pioneering work [10.1007/978-3-030-32239-7_88], Zhang et al. adopted the guided filter in the segmentation network and suggested incorporating the attention map $M$ into the energy minimization problem to estimate $a$ and $b$:
$\mathcal{E}(a_{k},b_{k})=\sum_{i\in\mathrm{w}_{k}}\left(M_{i}^{2}\left(a_{k}{I_{d}}_{i}+b_{k}-g_{i}\right)^{2}+\epsilon a_{k}^{2}\right).$ (6)
@@
In Eq. (6), the ${I_{d}}_{i}$ is the downsampled input feature from encoding parts, where $g_{i}$ denotes the gating signal from upward path at the decoding parts. Eq. (6) implies that an improved attention map can lead to better solutions. In the following, we are going to propose a method that can generate an improved attention map with full-scale information.
Network architecture
Figure 2: Network architecture of the proposed FSG-Net. In the feature representation network (left), down-convolution layers concatenate separately extracted features, connected via deep bottleneck structures to up-convolution layers. The red dashed boxes highlight feature merging from multiple scales. The guided convolution block (right) employs Guided Residual Modules (GRM) to refine multi-scale features. $S_{i}$ indicates each stage, and $y_{i}$ represents labels, where $y_{2}$ and $y_{3}$ are downsampled from the original label $y_{1}$ by interpolation.
@@
As can be seen in Fig. 2, the proposed network basically follows the U-Net architecture but has an additional guided convolution block after the feature representation network, unlike usual U-Net based structures. Contrary to the standard U-Net consisting of five stages, we have opted for four stages in FSG-Net based on the assumption that a wider receptive field is less critical for retinal vessel segmentation.
As illustrated on the left side of the FSG-Net architecture, the input to each down-convolution block, which is indicated by the bold green arrows in the figure, is augmented by concatenating features derived from an auxiliary convolutional path. These preserved features are subsequently forwarded to the corresponding up-convolution layers via skip connections, facilitating richer multi-scale feature reuse. The convolutional blocks within this stage adopt a deep bottleneck structure, which enhances the network’s representational capacity while maintaining computational efficiency. Furthermore, a lightweight spatial attention, inspired by the Convolutional Block Attention Module (CBAM) [Woo_2018_ECCV], is integrated into the bottleneck stage to refine spatial feature encoding with minimal parameter overhead.
@@
In the feature representation network, the key area to note is the feature merging section, indicated by the red dashed box. In this feature merging process, features from three paths—the current, upper, and lower stages—are concatenated and passed through a newly designed convolution block before being forwarded to the subsequent stages. This approach allows the integration and transmission of information across all scales. In this study, we refer to such integrated multi-scale information as full-scale information. The advantages of gathering the entire information in this manner will be further discussed in the Results section.
@@
In the right part of the model, namely the guided convolution block, compressed features reflecting full scale information are given as the current stage input to the guided residual module (GRM). Up-stage input from the higher stage is given as another input for guided filtering. After GRM, predictions are derived through convolution and activation. For training, a combination of the BCE loss and Dice loss with deep supervision was employed to improve segmentation performance of the edge-like structures [rs15061530]:
$L_{\text{DS}}=\sum_{d=1}^{D}\alpha_{d}\cdot\left(L_{\text{BCE}}^{(d)}+\lambda\cdot L_{\text{Dice}}^{(d)}\right),$ (7)
$\text{where}\quad L_{\text{BCE}}^{(d)}=-\frac{1}{S_{d}}\sum_{i=1}^{S_{d}}\left(y_{i,d}\log\hat{y}_{i,d}+(1-y_{i,d})\log(1-\hat{y}_{i,d}\right),$ (8)
$\text{and}\quad L_{\text{Dice}}^{(d)}=1-\frac{2\sum_{i=1}^{S_{d}}y_{i,d}\hat{y}_{i,d}+\epsilon}{\sum_{i=1}^{S_{d}}y_{i,d}+\sum_{i=1}^{S_{d}}\hat{y}_{i,d}+\epsilon}.$ (9)
In Eq. (7), $D$, $\alpha_{d}$ represent the number of prediction layers and weights for each prediction layer $d$, respectively. In Eq. (8) and Eq. (9), $S_{d}$ denotes the number of samples at the $d$-th prediction layer and $y_{i,d}$ and $\hat{y}_{i,d}$ represent target and prediction values at $i$-th index and at $d$-th prediction level, respectively.
@@
Modernized convolution block
The standard U-Net has a fundamental structure comprising skip connections and double convolution blocks. Recent studies have shown that incorporating different CNN structures can significantly improve the performance of the original model [liu2022convnet, HAN2022109512]. Inspired by the ConvNeXt [liu2022convnet], we have designed a convolutional block suitable for retinal vessel segmentation.
@@
Fig. 3(a) and (b) represent the depthwise residual block and inverted residual block, respectively. Fig. 3(c) shows the proposed convolution block. The structure of the block is an extension of the latest advancements in convolutional block development, characterized by features such as $\times1$ convolution, inverted bottleneck, and depth-wise convolution. Like other modernized convolution blocks, the proposed block employs a convolution with kernel size 2 and stride 2 in the first stage. By incorporating spatial and dimensional changes within the first block, we enable the design of a deep bottleneck stage in the down-convolution structure. This approach not only increases the information on the feature but also separates it from the bottleneck, allowing for more detailed feature representation. Furthermore, to maintain linearity, we utilize unique ReLU activation between each inverted residual block. For regularization purposes, we employ a learnable gamma parameter and apply a drop block before joining the identity block.
Guided convolution block
As shown in Fig. 2, the guided convolution block consists of GRM to refine the input from the feature representation network and a convolution to output the prediction map. Fig. 4 shows the overall process of GRM.
The purpose of GRM is to refine the feature from the feature representation network at the same stage. In GRM, both the features at the current stage and at the higher stage are used to generate an attention map, and then the attention map is multiplied with the current feature to generate the output, which is described in the right part of Fig. 4. Here we can expect an improved attention map $M$ can be estimated to solve Eq. (6) because the input feature from the feature representation network has full scale information. To further enhance the feature refinement, a residual block is introduced after the attention-guided filtering, as shown in the left part of Fig. 4. By passing the features through the residual block, a stable map is generated for both the deep supervision and subsequent layers. Moreover, we incorporate a $1\times 1$ convolution to preserve the semantic information present in the feature maps.
@@
Experiments
Training Techniques
In retinal vessel segmentation, the performance of studies is often determined by subtle gaps. We believe that these subtle gaps are highly influenced by the choice of hyperparameters and training/inference environment. To address this imbalance, we fixed all hyperparameters for training and inference. We empirically found that RandAugment [Cubuk_2020_CVPR_Workshops] with a specific scale did not work well on medical datasets; therefore, we customized it to better suit our datasets. Training techniques include blur, color jitter, horizontal flip, perspective transformation, resize, crop, and CutMix [Yun_2019_ICCV].
Implementation Details
[…]
@@
lr scheduler Linear warm-up, Cosine annealing
@@
[…]
random jitter b=0.2, c=0.2, s=0.2, h=0.1, prob=0.8
@@
random perspective s=0.3, prob=0.3
random random resize s=[0.5, 2.0], prob=0.8
@@
Datasets
[…]
Table 2: Comparison of segmentation performance for CNN-based networks.
Architecture
mIoU
F1 score
Acc
AUC
Sen
MCC
Rank Avg
@@
U-Net 83.857 82.956 97.013 97.853 83.449 81.456 5.7
U-Net++ 81.228 79.564 96.524 96.271 77.802 77.830 14.0
U-Net3+ Deep 83.909 83.030 97.017 98.082 83.721 81.520 3.5
ResU-Net 83.862 82.953 97.021 97.766 83.226 81.453 6.5
ResU-Net++ 83.729 82.783 97.001 97.708 82.791 81.263 10.2
SAU-Net 83.368 82.334 96.925 97.616 82.311 80.782 12.2
DCASU-Net 83.743 82.808 96.996 97.838 83.080 81.290 8.7
AG-Net 83.176 82.111 96.882 97.628 82.155 80.540 12.8
AttU-Net 83.958 83.080 97.039 97.844 83.422 81.584 3.5
R2U-Net 83.555 82.580 96.952 97.879 82.961 81.038 9.7
ConvU-NeXt 83.800 82.882 97.012 97.835 83.019 81.367 7.8
FR-UNet 83.884 82.995 97.007 98.158 83.869 81.485 4.3
HRNet 83.938 83.829 97.325 97.860 82.963 81.506 5.0
FSG-Net (ours) 84.068 83.229 97.042 98.235 84.207 81.731 1.2
@@
U-Net 85.924 84.873 97.754 98.341 84.361 83.713 3.7
U-Net++ 81.514 79.061 97.022 94.479 75.539 77.764 13.8
U-Net3+ Deep 85.824 84.829 97.707 99.146 85.522 83.626 3.7
ResU-Net 85.964 84.872 97.767 98.050 83.997 83.726 4.0
ResU-Net++ 83.185 81.358 97.319 95.221 78.456 80.196 12.7
SAU-Net 85.158 84.061 97.604 97.784 84.015 82.885 8.0
DCASU-Net 84.423 83.064 97.454 98.362 83.583 81.771 9.7
AG-Net 84.811 83.766 97.565 98.403 83.347 82.516 8.5
AttU-Net 85.848 84.772 97.694 99.050 86.226 83.588 4.0
R2U-Net 83.727 81.786 97.468 93.457 77.048 80.810 12.3
ConvU-NeXt 85.339 84.186 97.658 97.866 83.401 82.998 7.5
FR-UNet 84.815 83.496 97.577 96.637 81.872 82.327 10.2
HRNet 85.555 84.165 97.982 98.190 84.604 83.197 5.2
FSG-Net (ours) 86.118 85.100 97.746 98.967 86.608 83.958 1.8
@@
U-Net 82.065 80.159 97.404 99.368 85.370 79.002 7.0
U-Net++ 81.512 79.415 97.321 99.362 84.143 78.201 9.2
U-Net3+ Deep 82.489 80.697 97.483 99.506 85.740 79.558 3.7
ResU-Net 81.104 78.810 97.320 99.278 81.311 77.479 12.5
ResU-Net++ 73.276 67.411 95.966 96.223 67.782 65.593 14.0
SAU-Net 81.335 79.100 97.338 99.407 82.265 77.802 9.8
DCASU-Net 82.254 80.368 97.483 99.329 83.916 79.161 6.3
AG-Net 82.158 80.272 97.440 99.544 84.817 79.070 5.2
AttU-Net 82.562 80.742 97.546 99.430 83.907 79.537 4.3
R2U-Net 81.250 78.944 97.384 98.792 80.054 77.625 11.8
ConvU-NeXt 81.752 79.704 97.405 99.402 83.009 78.439 8
FR-UNet 81.330 79.170 97.269 99.544 84.744 77.944 8.7
HRNet 82.849 81.021 97.744 99.650 85.750 79.987 1.2
FSG-Net (ours) 82.680 81.019 97.515 99.378 85.995 79.899 3.0